OpenStudy (anonymous):

Sang-Jin has only quarters and nickels in her piggy bank. She has 66 coins in the bank with a value of $11.50. How many quarters are in Sang-Jin’s piggy bank? _________

OpenStudy (anonymous):

@Loser66 @wio

OpenStudy (dtan5457):

Your system of equations should be 0.25x+0.05x=11.50 x+y=66 I'd multiply the first equation by 100 to get rid of the decimals.

OpenStudy (dtan5457):

Know how to solve from here?

OpenStudy (anonymous):

no, sorry. you think you could explain a bit more? :) @dtan5457

OpenStudy (dtan5457):

(0.25x+0.05x=11.50)100?

OpenStudy (anonymous):

x=3833.33333333? @dtan5457

OpenStudy (dtan5457):

You don't solve for x like that. Once you multiply by 100 25x+5y=1150 the 2nd one is still x+y=66 , which I would multiply by 25 25x+25y=1650 25x+5y=1150 subtract 20y=500 y=? (this is for nickels)

OpenStudy (anonymous):

25?

OpenStudy (dtan5457):

Exactly. So there is 25 nickels, 66 coins in total, how many quarters?

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

66-25=?

OpenStudy (anonymous):

41. Sorry, I typed the wrong number into my calculator. :)

OpenStudy (dtan5457):

Yeah. So 25 nickels 41 quarters